A rectangular table is 45 inches wide and 64 inches long.A table cloth hangs over the 10 inches at each edge.What is the area of the table cloth.

Answer by stanbon(75887)

A rectangular table is 45 inches wide and 64 inches long.A table cloth hangs over the 10 inches at each edge.What is the area of the table cloth.
----
width = 45 + 10 + 10 = 65
length = 64 + 10 + 10 = 84
------
Area = 65*84 sq. in
-----------
Cheers,
Stan H.
-------------------